OpenStudy (anonymous):

how can u do this :(6x+7) (3x+2)

OpenStudy (anonymous):

use foil 18x^2+33x+14

OpenStudy (anonymous):

Use the FOIL method. (6x+7) (3x+2) F = 6x * 3x (18x^2) O = 6x * 2 (12x) I = 7 * 3x (21x) L = 7 * 2 (14) now put it all together 18^2 + 12x + 21x + 14 Now simplify by combining like terms 18^2 +33x + 14 F = First pair of numbers O = Outer pair of numbers I = Inner pair of numbers L = last pair of numbers